# tidychangepoint: a unified framework for analyzing changepoint detection in univariate time series

### v2

@@ 1.5 Our contribution
In this paper, we present several substantive improvements to the existing state-of-the-art. First, the tidychangepoint package (Baumer et al. 2024) provides a unified interface that makes working with existing changepoint detection algorithms in a tidyverse-compliant syntax easy (Section 3). This reduces friction for anyone who wants to compare the results of algorithms, models, or penalized objective functions across different packages. Second, the architecture of the package is easily extended to incorporate new changepoint detection algorithms (Section 4), new parametric models (Section 5), and new penalized objective functions (Section 6). Indeed, our careful separation of these three essential elements of changepoint detection is in and of itself clarifying. Third, to the best of our knowledge, tidychangepoint is the first R package that allows users to run genetic algorithms for changepoint detection without having to write custom code. With GA as a backend, tidychangepoint provides a user-friendly interface for mixing-and-matching various parametric models, penalized objective functions, and initial populations in ways that previously required non-trivial programming by the user. Since tidychangepoint wraps functionality from other packages it adds only modest computational overhead, and delivers correct results (Section 7). Finally, tidychangepoint includes new or updated data sets packaged as time series objects about particulate matter in Bogotá, monthly rainfall in Medellín, temperature in Central England, and batting performance in Major League Baseball. We illustrate the utility of the package in Section 2 and conclude in Section 8.
@@ 2.1 Temperatures in Central England
In Section 1.1, we noted that the comparative analysis across different models presented in Shi et al. (2022) required writing custom code for each model fit. In Section 5, we list the many parametric models that tidychangepoint includes. These include the “meanshift” and “trendshift” models, either of which can include AR(1) lagged errors, and either of which can be fit using the BIC or MDL penalties (see Section 6).
@@
Next, using functionality from the purrr package (Wickham and Henry 2023), we simply iterate the segment() function—which is the main workhorse in tidychangepoint—over these combinations. In each case we employ a genetic algorithm that can produce up to 5000 generations (maxiter), although it will stop if there is no improvement in the fitness after 100 consecutive generations (run).
@@
In this case, the trendshift model with white noise errors produces the best fitness score with the BIC penalty, while adding white noise errors produces the best fitness score with the MDL penalty. These results accord with the findings by Shi et al. (2022).
@@ 2.2 The designated hitter in Major League Baseball
Sometimes we know the truth about when changepoints occur. For example, Figure 4 shows the difference in batting average between the American and National Leagues of Major League Baseball (MLB) from 1925–2020. The introduction of the designated hitter rule in the American League in 1973 led to noticeable differences in batting performance relative to the National League (which did not adapt the rule until 2020), because the designated hitter employed by American League teams was always a better hitter than the pitcher (who was forced to bat in the National League). Note that in Figure 4, the American League had a higher batting average in every single year from 1973 to 2020. Thus, the year 1973 is a known changepoint, because the data generating process changed before and after that year.
Figure 4: Difference in the rate of home runs per plate appearance between the American and National Leagues in Major League Baseball, 1925–2022. The designated hitter rule was adopted in 1973 by the American League, but not until 2020 by the National League. Note how, during the period from 1973 to 2020 when only the American League employed a designated hitter, the difference in batting average was always negative.
@@
    mlb_bavg, method = "ga"model_fn = fit_meanshift_norm,
    penalty_fn = MDL, maxiter = 500run = 50
Because each of the resulting objects are of class tidycpt (see Section 3), comparing the results across the four algorithms is easy using map() (from the purrr package (Wickham and Henry 2023)). First, we examine the changepoint sets.
@@
  map(changepoints, use_labels = TRUE|>
@@
To compare other aspects of the fits, we can iterate the glance() function from the broom package (Robinson, Hayes, and Couch 2024) over that same list. Since glance() always returns a one-row tibble, we can then combine those tibbles into one using the list_rbind() function. Care must be taken to ensure that these comparisons are meaningful. While comparing, say, the elapsed time across these algorithms and models is fair, comparing the fitness values is more fraught. MDL scores are not comparable to BIC or MBIC scores, for example.
@@
3 wbs      1.4.1   Wild Bin~ <list [1]> meanshift~ MBIC          0   0.010 secs
@@
Figure 5: Comparison of changepoint sets returned by four algorithms. Which provides the best fit to the data?
Visually, we could use a rug plot (see ggplot2::geom_rug()) to show the distribution of changepoints across algorithms (see Figure 6). Three of the four algorithms found a changepoint near the true changepoint of 1973. The other changepoints are false positives with respect to the designated hitter rule, but may very well reflect other changes in batting performance that we simply haven’t been thinking about. We suspect that the recent decrease in the difference in batting average could reflect fewer plate appearances taken by pitchers due to the increased use of relief pitchers.
@@ 2.2.1 Investigating algorithmic performance
As expected, the MBIC value of the meanvar model with the true changepoint set is worse (further from $\infty$- ∞).
@@
list(49changepoints(mlb_mods[["pelt_meanvar"]])) |>
@@ 2.3 Particular matter in Bogotá, Colombia
    method = "ga-coen"maxiter = 1000run = 100,
@@
Alternatively, a careful look at the data suggest that the trendshift model, which allows for a linear slope within each region, might be more appropriate. Setting the method argument of segment() to ga allows us to employ a custom genetic algorithm for which we can specify a model via the model_fn argument. Here, we use the MDL() penalty function and the log_gabin_population() function. While Figure 8 picks up on the COVID-19-related changepoint, Figure 8 shows that the trendshift model also fits the data very well.
@@
    method = "ga"model_fn = fit_trendshift, penalty_fn = MDL,
@@
(b) A genetic algorithm with a trendshift model.
Figure 8: Changepoint sets for two genetic algorithms applied to the Bogotá particulate matter data. Both pick up on the COVID-19-related changepoint in March of 2020. The latter includes linear trends in each region.
@@
(b) The diagnose() function applied to a model object of class mod_cpt (a genetic algorithm with a trendshift model).
@@ 3 Architecture of tidychangepoint
  segment(method = "pelt"model_fn = fit_meanvar, penalty = "BIC")
@@ 3.2 Methods available
tidycpt objects implement methods for the generic functions as.model(), as.segmenter(), as.ts(), changepoints(), diagnose(), fitness(), model_name(), regions, and plot(), as well as three generic functions from the broom package: augment(), tidy(), and glance().
@@
[11] print        regions      tidy
@@
For the most part, methods for tidycpt objects typically defer to the methods defined for either the segmenter or the model, depending on the user’s likely intention. To that end, both segmenters and models implement methods for the generic functions as.ts(), changepoints(), model_name(), and nobs().
@@
The changepoints() function return the indices of the changepoint set, unless the use_labels argument is set to TRUE.
@@
changepoints(mlb_pelt_meanvar_bic, use_labels = TRUE|>
@@
  • tidy() returns a tibble (Müller and Wickham 2023) that provides summary statistics for each region. These include any parameters that were fit, which are prefixed in the output by param_.
@@
1 changep~ 2.3     PELT      <list [1]> meanvar    BIC        -782. 0.009 secs
@@ 4.1 Segmenting algorithms provided
  • A variety of genetic algorithms as implemented by GA (Scrucca 2024). Specific variants include Shi’s algorithm (Shi et al. 2022) and Coen’s algorithm (Suárez-Sierra, Coen, and Taimal 2023).
@@ 5 Models
[7] "regions"   "residuals" "tidy"
@@
1 tidy~ 1.0.0.~ meanvar   <NULL>        4 0.00507   367. -706. -671. -689. -674.
@@ 5.1 Model-fitting functions provided
fit_trendshift() $32m+$ 3 $\sigma^{2},\mu_{0},\ldots,\mu_{m},\beta_{0},\ldots,\beta_{m}$ special case of fit_lmshift() with $p=1$
fit_trendshift_ar1() $42m+$ 4 $\sigma^{2},\phi,\mu_{0},\ldots,\mu_{m},\beta_{0},\ldots,\beta_{m}$
@@
fit_meanvar() $22m+$ 2 $\sigma_{0}^{2},\ldots,\sigma_{m}^{2},\mu_{0},\ldots,\mu_{m}$
fit_nhpp() $22m+$ 2 $\alpha_{0},\ldots,\alpha_{m},\beta_{0},\ldots,\beta_{m}$
@@ 5.3 Spotlight: Non-homogenous Poisson process model
$\ln{g(\theta_{\tau}|y)}\propto\ln{L_{NHPP}(y|\theta_{\tau})}+\ln{g(\theta_{\tau})}\,.$
@@ 6 Penalized objective functions
$f(\tau,M(y|\theta_{\tau}))=P_{f}(\tau,n)-2\ln{L_{M}(y|\theta_{\tau})}\,,$
where $P_{f}$ is a function that guards against overfitting by penalizing large changepoint sets—that is, it is a monotonically increasing function of $m$. The stats package (R Core Team 2024) provides generic methods for the well-known penalty functions AIC() and BIC(), which work seamlessly on tidychangepoint models because they all implement methods for logLik(). In addition, tidychangepoint provides generic functions and methods for four additional penalty functions: HQC(), MBIC(), MDL(), and BMDL(). For compatibility with changepoint we also provide the alias SIC() for BIC(). tidychangepoint supports all of the penalty functions supported by changepoint, with the exception of the “Asymptotic” and “CROPS” penalties, which are not fully generalizable.
All penalty functions return 0 for the special case in which $m=0$. For ease of explanation, let the vector $\ell_{j}=\tau_{j+1}-\tau_{j}$ for $11\leq j\leq m+$ 1 encode the lengths of the $m+1$ regions defined by the changepoint set $\tau$.
@@ 6.1.2 Modified Bayesian Information Criterion
Following Zhang and Siegmund (2007) and Li and Lund (2012), we define the MBIC as:
@@ 6.1.3 Minimum Descriptive Length
As described in Shi et al. (2022) and Li and Lund (2012), we define the MDL as:
$P_{MDL}(\tau,n)=\frac{a(\theta_{\tau})}{2}\cdot\sum_{j=0}^{m}\log{\ell_{j}}+2\ln{m}+\sum_{j=2}^{m}\ln{\tau_{j}}+\left(2+b(\theta_{\tau})\right)\ln{n}\,.$
@@ 7.1 Correctness
  segment(method = "pelt"penalty = "BIC")

### v3

@@ 1.7 Our contribution
In this paper, we present several substantive improvements to the existing state-of-the-art. First, the tidychangepoint package (Baumer et al. 2025) provides a unified interface that makes working with existing changepoint detection algorithms in a tidyverse-compliant syntax easy (Section 3). This reduces friction for anyone who wants to compare the results of algorithms, models, or penalized objective functions across different packages. Second, the architecture of the package is easily extended to incorporate new changepoint detection algorithms (Section 4), new parametric models (Section 5), and new penalized objective functions (Section 6). Indeed, our careful separation of these three essential elements of changepoint detection is in and of itself clarifying. Third, to the best of our knowledge, tidychangepoint is the first R package that allows users to run genetic algorithms for changepoint detection without having to write custom code. With GA as a backend, tidychangepoint provides a user-friendly interface for mixing-and-matching various parametric models, penalized objective functions, and initial populations in ways that previously required non-trivial programming by the user. Since tidychangepoint wraps functionality from other packages it adds only modest computational overhead, and delivers correct results (Section 7). Finally, tidychangepoint includes new or updated data sets packaged as time series objects about particulate matter in Bogotá, monthly rainfall in Medellín, temperature in Central England, and batting performance in Major League Baseball. We illustrate the utility of the package in Section 2 and conclude in Section 8.
@@ 2.1 Temperatures in Central England
In Section 1.3, we noted that the comparative analysis across different models presented in Shi et al. (2022) required writing custom code for each model fit. In Section 5, we list the many parametric models that tidychangepoint includes. These include the “meanshift” (i.e., piecewise-constant) and “trendshift” (i.e., piecewise-linear) models, either of which can include AR(1) lagged errors, and either of which can be fit using the BIC or MDL penalties (see Section 6).
@@
Next, using functionality444purrr::map2() is similar to base::mapply(). from the purrr package (Wickham and Henry 2025), we simply iterate the segment() function—which is the main workhorse in tidychangepoint—over these combinations. In each case we employ a genetic algorithm that can produce up to 5000 generations (maxiter), although it will stop if there is no improvement in the fitness (i.e., the value of the penalized objective function) after 100 consecutive generations (run).
@@
In this case, the trendshift model with white noise errors produces the best fitness score with the BIC penalty, while adding AR(1) errors produces the best fitness score with the MDL penalty. These results accord with the findings by Shi et al. (2022).
@@ 2.2 The designated hitter in Major League Baseball
Sometimes we know the truth about when changepoints occur. For example, Figure 3 shows the difference in batting average between the American and National Leagues of Major League Baseball (MLB) from 1925–2019. The introduction of the designated hitter rule in the American League in 1973 led to noticeable differences in batting performance relative to the National League (which did not adapt the rule until 2020), because the rule allowed teams in the American League to designate another player to hit in place of the pitcher, whereas in the National League, the current pitcher continued to be forced to hit. Most pitchers are poor hitters relative to a Major League non-pitcher, and so the designated hitter rule allowed American League teams to bat 9 non-pitchers all the time, whereas National League teams generally had 8 non-pitchers and one pitcher in the batting lineup. The goal of the rule change was to increase offensive performance in the American League, and it worked. Note that in Figure 3, the American League had a higher batting average in every single year from 1973 to 2019. Thus, the year 1973 is a known changepoint, because the data generating process changed before and after that year.
Figure 3: Difference in the rate of home runs per plate appearance between the American and National Leagues in Major League Baseball, 1925–2019. The designated hitter rule was adopted in 1973 by the American League, but not until 2020 by the National League. Note how, during the period from 1973 to 2019 when only the American League employed a designated hitter, the difference in batting average was always negative.
@@
mlb_bavg, method = "ga", model_fn = fit_meanshift_norm,
penalty_fn = MDL, maxiter = 500, run = 50
Because each of the resulting objects are of class tidycpt (see Section 3), comparing the results across the four algorithms is easy using map()555purrr::map() is similar to base::lapply(). (from the purrr package (Wickham and Henry 2025)). First, we examine the changepoint sets.
@@
map(changepoints, use_labels = TRUE) |>
@@
To compare other aspects of the fits, we can iterate the glance() function from the broom package (Robinson et al. 2025) over that same list. Since glance() always returns a one-row tibble, we can then combine those tibbles into one using the list_rbind() function. Care must be taken to ensure that these comparisons are meaningful. While comparing, say, the elapsed time across these algorithms and models is fair, comparing the fitness values is more fraught. MDL scores are not comparable to BIC or MBIC scores, for example.
@@
3 wbs      1.4.1   Wild Bin~ <list [1]> meanshift~ MBIC       -473.  0.010 secs
@@
Figure 4: Comparison of changepoint sets returned by four algorithms. The question of which provides the best fit is a matter of interpretation.
Visually, we could use a rug plot (see ggplot2::geom_rug()) to show the distribution of changepoints across algorithms (see Figure 5). Three of the four algorithms found a changepoint near the true changepoint of 1973. The other changepoints are false positives with respect to the designated hitter rule, but may very well reflect other changes in batting performance that we simply haven’t been thinking about. For example, we hypothesize that the recent decrease in the difference in batting average across the two leagues could be a result of the increased use of relief pitchers (who are almost always replaced by a pinch hitter when their turn at bat comes up) at the expense of starting pitchers (who can’t be pinch-hit for because they have to stay in the game)—a long-term change in pitching strategy. If this trend continued to its extreme and pitchers never batted, the designated hitter rule would be moot and the American League’s advantage would evaporate.
@@ 2.2.1 Investigating algorithmic performance
As expected, the MBIC value of the meanvar model with the true changepoint set is worse (further from -$\infty$).
@@
list(49, changepoints(mlb_mods[["pelt_meanvar"]])) |>
@@ 2.3 Particular matter in Bogotá, Colombia
method = "ga-coen", maxiter = 1000, run = 100,
@@
A more thorough analysis might use the decompose() or stl() functions from the stats package to unravel the seasonality of the particulate matter data. Instead, for the purposes of simplicity here we simply experiment with different models. Setting the method argument of segment() to ga allows us to employ a custom genetic algorithm for which we can specify a model via the model_fn argument. In this case, we use the fit_lmshift() function (see Section 5.1) to fit a third-degree polynomial (note the model_fn_args argument) to each region. Here, we use the MDL() penalty function and the log_gabin_population() function. Both models in Figure 7 pick up on the COVID-19-related changepoint. The piecewise constant model in Figure 7 ignores the seasonality present in the data, while the Figure 7 is undoubtedly overfit.
@@
method = "ga", model_fn = fit_lmshift,
model_fn_args = list(deg_poly = 3),
penalty_fn = MDL,
@@
(b) A genetic algorithm with a polynomial trend model.
Figure 7: Changepoint sets for two genetic algorithms applied to the Bogotá particulate matter data. Both pick up on the COVID-19-related changepoint in March of 2020. The latter includes 3rd degree polynomial trends in each region.
@@
(b) The diagnose() function applied to a model object of class mod_cpt (a genetic algorithm with a polynomial trend model).
@@ 3 Architecture of tidychangepoint
segment(method = "pelt", model_fn = fit_meanvar, penalty = "BIC")
@@ 3.2 Methods available
tidycpt objects implement methods for the generic functions as.model(), as.segmenter(), as.ts(), changepoints(), diagnose(), fitness(), model_name(), regions(), plot(), and summary(), as well as three generic functions from the broom package: augment(), tidy(), and glance().
@@
[11] print        regions      summary      tidy
@@
For the most part, methods for tidycpt objects typically defer to the methods defined for either the segmenter or the model, depending on the user’s likely intention. To that end, both segmenters and models implement methods for the generic functions as.ts(), changepoints(), model_name(), nobs(), plot(), and print().
@@
The changepoints() function returns the time indices of the changepoint set, unless the use_labels argument is set to TRUE.
@@
changepoints(mlb_pelt_meanvar_bic, use_labels = TRUE) |>
@@
  • tidy() returns a tibble (Müller and Wickham 2025) that provides summary statistics for each region. These include any parameters that were fit, which are prefixed in the output by param_.
@@
1 changep~ 2.3     PELT      <list [1]> meanvar    BIC        -782. 0.01 secs
@@ 4.1 Segmenting algorithms provided
  • Segmented regression models as implemented by segmented (V. M. R. Muggeo 2025)
  • A variety of genetic algorithms as implemented by GA (Scrucca 2024) and changepointGA (M. Li and Lu 2025). Specific variants include Shi’s algorithm (Shi et al. 2022) and Coen’s algorithm (Suárez-Sierra, Coen, and Taimal 2023).
@@
cptga changepointGA tidycptga segment_cptga() changepointGA::cptga()
@@ 5 Models
 [7] "regions"   "residuals" "summary"   "tidy"
@@
1 tidy~ 1.0.3   meanvar   <NULL>        4 0.00507   367. -706. -671. -689. -674.
@@ 5.1 Model-fitting functions provided
fit_trendshift() $2m+3$ $\sigma^{2},\mu_{0},\ldots,\mu_{m},\beta_{0},\ldots,\beta_{m}$ special case of fit_lmshift() with $p=1$
fit_trendshift_ar1() $2m+4$ $\sigma^{2},\phi,\mu_{0},\ldots,\mu_{m},\beta_{0},\ldots,\beta_{m}$
@@
fit_meanvar() $2m+2$ $\sigma_{0}^{2},\ldots,\sigma_{m}^{2},\mu_{0},\ldots,\mu_{m}$
fit_nhpp() $2m+2$ $\alpha_{0},\ldots,\alpha_{m},\beta_{0},\ldots,\beta_{m}$
@@ 5.3 Spotlight: Non-homogenous Poisson process model
$\ln{g(\theta_{\tau}|y)}\propto\ln{L_{NHPP}(y|\theta_{\tau})}+\ln{g(\theta_{\tau})}\,$,
@@ 6 Penalized objective functions
$f(\tau,M(y|\theta_{\tau}))=\underbrace{P_{f}(\tau,n)}_{penalty}-2\ln{\underbrace{L_{M}(y|\theta_{\tau})}_{likelihood}}\,$,
where $P_{f}$ is a function that guards against overfitting by penalizing large changepoint sets—that is, it is a monotonically increasing function of $m$ (the number of changepoints). Recall that $\tau$ is the set of changepoints. The stats package (R Core Team 2025) provides generic methods for the well-known penalty functions AIC() and BIC(), which work seamlessly on tidychangepoint models because they all implement methods for logLik(). In addition, tidychangepoint provides generic functions and methods for four additional penalty functions: HQC(), MBIC(), MDL(), and BMDL(). For compatibility with changepoint we also provide the alias SIC() for BIC(). tidychangepoint supports all of the penalty functions supported by changepoint, with the exception of the “Asymptotic” and “CROPS” penalties, which are not fully generalizable.
All penalty functions return 0 for the special case in which $m=0$. For ease of explanation, let the vector $\ell_{j}=\tau_{j+1}-\tau_{j}$ for $0\leq j\leq m$ encode the lengths of the $m+1$ regions defined by the changepoint set $\tau$.
@@ 6.1.2 Modified Bayesian Information Criterion
Following Zhang and Siegmund (2007) and S. Li and Lund (2012), we define the MBIC as:
@@ 6.1.3 Minimum Descriptive Length
As described in Shi et al. (2022) and S. Li and Lund (2012), we define the MDL as:
$P_{MDL}(\tau,n)=\frac{a(\theta_{\tau})}{2}\cdot\sum_{j=0}^{m}\log{\ell_{j}}+2\ln{m}+\sum_{j=2}^{m}\ln{\tau_{j}}+\left(2+b(\theta_{\tau})\right)\ln{n}\,$,
@@ 7.1 Correctness
segment(method = "pelt", penalty = "BIC")